# Short-term CO2 emissions forecasting: insight from the Italian electricity market

### v1

@@ Abstract
This study investigates the short-term forecasting of carbon emissions from electricity generation in the Italian power market. Using hourly data from 2021 to 2023, several statistical models and forecast combination methods are evaluated and compared at the national and zonal levels. Four main model classes are considered: (i) linear parametric models, such as seasonal autoregressive integrated moving average and its exogenous-variable extension; (ii) functional parametric models, including seasonal functional autoregressive models, with and without exogenous variables; (iii) (semi) non-parametric and possibly non-linear models, notably the generalised additive model (GAM) and TBATS (trigonometric seasonality, Box-Cox transformation, ARMA errors, trend, and seasonality); and (iv) a semi-functional approach based on the K-nearest neighbours. Forecast combinations are also considered including simple averaging, the optimal Bates and Granger weighting scheme, and a selection-based strategy that chooses the best model for each hour. The overall findings indicate that GAM reports the most accurate forecasts during the daytime hours, while functional parametric models perform best during the early morning period. GAM can also be considered the best individual model according to the hourly average root mean square error and the Diebold-Mariano (DM) test. Among the combination methods, the selection-based approach consistently outperforms all individual models and forecast combinations, resulting in a substantial reduction in the root mean square error compared to single models and a primary choice for the DM test. These findings underline the value of hybrid forecasting frameworks in improving the accuracy and reliability of short-term carbon emissions predictions in power systems.
@@ 1 Introduction
Carbon dioxide (CO2) emissions are a major driver of global climate change. According to the emissions database for global atmospheric research (European Commission,, 2024), in 2023, China, the world’s largest emitter, accounted for 33.98% of global CO2 emissions, followed by the USA (12%), India (7.57%), Europe (6.44%) and Russia (5.30%). The power sector remains a key contributor, responsible for 38.24% of total CO2 emissions in the same year. In particular, the power sectors of China, the USA, India, Europe and Russia contributed 16.59%, 3.75%, 3.53%, 2.26%, and 1.64% of global CO2 emissions, respectively. As a result, these countries represented 72.61% of global CO2 emissions from the power sector.
@@
The amount of CO2 emissions from electricity generation depends on the share of renewable sources (e.g. hydro, solar, and wind) and nonrenewable sources (e.g. coal, oil, and natural gas). The mix of sources changes over time according to the mechanisms of the electricity market and weather conditions (Kocak et al.,, 2023). Although renewable energy generation reached a global record of 927 TWh in 2024, CO2 emissions from the power sector increased by 1.6% compared to 2023. This increase was mainly driven by an increase in global electricity demand, particularly due to extreme heat in countries such as India (Carbon Brief,, 2025). The rise in fossil fuel use highlights the ongoing challenge of decarbonising under climate stress.
@@
Over time, most countries have made efforts to reduce CO2 emissions through international initiatives. The Paris Agreement, signed in 2015, aims to keep global warming well below $2^{\circ}$ above pre-industrial levels, with the aim of staying below $1.5^{\circ}$ (United Nations, 2015a, ). It also sets specific targets for the power sector, including a 20% reduction in CO2 emissions, a 20% increase in the use of renewable energy, and a 20% improvement in energy efficiency worldwide. These actions also support Sustainable Development Goal 7 (SDG 7), introduced in the same year with the 2030 Agenda for Sustainable Development, which focusses on ensuring access to affordable, reliable, and clean energy for all (United Nations, 2015b, ). Making the power sector more sustainable is essential to reduce CO2 emissions and address climate change.
@@
As part of the global shift towards cleaner energy in line with the climate targets of the United Nations (United Nations,, 2020), more attention is paid to strategies that reduce CO2 emissions from electricity generation.
@@
From the supply side, proposed solutions include the integration of renewable energy sources with energy storage systems to manage fluctuations in renewable electricity generation. Batteries, fuel cells, or hydro reservoirs can be programmed to charge during periods of high renewable generation, typically associated with low emissions, and discharge when renewable output is low and consequently emissions are higher (Haddadian et al.,, 2015; International Energy Agency,, 2019; He et al.,, 2024).
@@
On the demand side, there is growing interest in strategies that shift flexible energy demand from periods of high emissions to periods of low emissions. Smart charging represents an opportunity to reduce CO2 emissions by scheduling charging sessions during periods of lower emissions in electricity generation (Jochem et al.,, 2015). For example, residential charging of electric vehicles can be strategically planned when CO2 emissions are low (Huber and Weinhardt,, 2018; Huber et al.,, 2021). Even if electric vehicle drivers are not experts in the electric system (Biresselioglu et al.,, 2018), their willingness to adopt flexible charging, such as accepting delays, increases when it contributes to greater integration of renewable energy and reduces charging costs and carbon emissions (Will and Schuller,, 2016; Jung et al.,, 2019).
@@
Consumers typically schedule electricity usage based on price forecasts to minimise costs (Voronin and Partanen,, 2013). However, price-based scheduling can lead to lower emissions only when renewable energy production exceeds the needed (Leerbeck et al.,, 2020). Otherwise, it may result in a phenomenon known in the literature as the “merit order emission dilemma”, where low-cost but high-emission sources, such as coal, are preferred (Wagner et al.,, 2002; Fleschutz et al.,, 2021). Recently, more attention has been paid to the environmental impact of electricity consumption, with the aim of reducing CO2 emissions by promoting a shift to lower-carbon energy sources. In this context, electricity usage could be scheduled not only based on forecasted prices but also based on predicted CO2 emissions, as a primary or secondary objective (Bokde et al.,, 2021). Therefore, accurate CO2 forecasts are essential to support emission-based scheduling of flexible demand and storage, at least until electricity generation becomes fully renewable.
@@ 1.1 Literature review
In the CO2 forecasting framework, a distinction can be made between short-term and long-term CO2 emissions forecasting. Long-term forecasting, e.g. annual projections, enables policymakers to plan and manage future carbon footprints. By contrast, short-term CO2 emissions forecasting is essential for scheduling electricity production and consumption in ways that minimise resulting emissions over a day-ahead horizon. Although long-term forecasting has received substantial attention in the literature, studies focused on short-term forecasting are limited (Bokde et al.,, 2021; Leerbeck et al.,, 2020). Finenko and Cheah, (2016) analyse hourly CO2 emissions from electricity generation in Singapore, highlighting hourly patterns due to both supply-side factors (e.g., renewable integration) and demand-side dynamics (e.g., peak vs. off-peak loads). Mason et al., (2018) applied evolutionary neural networks to forecast short-term CO2 emissions in Ireland over a multi-hour horizon. Lowry, (2018) forecasted the CO2 intensity of the United Kingdom’s electricity grid on a day-ahead horizon using seasonal autoregressive integrated moving average (SARIMA) models and artificial neural networks. Leerbeck et al., (2020) propose a machine learning algorithm to forecast short-term average and marginal CO2 emissions intensities in the Danish bidding zone DK2. The proposed algorithm combines multiple linear regression models through Softmax-weighted averaging, with residuals corrected using a SARIMA model with exogenous variables (SARIMAX). Bokde et al., (2021) introduced two time series decomposition methods to forecast short-term CO2 emissions from electricity generation over the 48-hour horizon. The proposed methods are also compared with SARIMA models, model-based on differenced pattern sequences, feed-forward neural networks, pattern sequence-based forecasting methods, and support vector machine. Short-term CO2 emissions forecasts are obtained for five European countries: France, Germany, Poland, Denmark, and Norway. Maji et al., 2022b introduced the day-ahead carbon forecasting system, which forecasts the electricity generation of each source through an artificial neural network, then integrates these forecasts with the emission rates associated with each source to calculate the overall carbon emissions and carbon intensity. Maji et al., 2022a introduced CarbonCast, a machine learning system to forecast multi-day carbon intensities. CarbonCast has a hierarchical design that combine neural networks, weather forecasts, and historical data. Many of these studies (Lowry,, 2018; Leerbeck et al.,, 2020; Maji et al., 2022b, ; ElectricityMap,, 2022) provide forecasts for the day ahead, while some others provide multi-day forecasts (Mason et al.,, 2018; Bokde et al.,, 2021; Maji et al., 2022a, ).
@@
Jin et al., (2024) provided a comprehensive review of carbon emission prediction models, identifying four main categories [^1] This classification closely aligns with that identified by Damon and Guillas, (2002), who grouped models into three main categories: (i) time series and regression models-including ARMA models, threshold autoregressive models, generalized autoregressive conditional heteroskedasticity (GARCH) models, non-parametric regression models, generalized additive models, and adaptive regression splines; (ii) neural networks; and (iii) classification and regression tree models., [^2] Another interesting literature review on carbon emission forecast and model classification is provided by Tian et al., (2025).: statistical models, shallow intelligent models, neural networks, and combined models.
@@
Statistical models are the most used because of their interpretability and reliance on historical data. Among them, the grey model is particularly popular for scenarios with limited data (Deng,, 1982). Regression models, such as linear and panel regression, have been used to analyse trade-related emissions and regional trends (Singh and Kennedy,, 2015). Furthermore, time series models such as SARIMA and autoregressive distributed lag models (Lin and Agyeman,, 2019) have proven effective in capturing temporal dynamics, offering good forecasting performance in various applications (Li et al.,, 2018; Malik et al.,, 2020; Kaur et al.,, 2023). Despite their strengths in modelling trends and seasonality, these models may struggle with abrupt fluctuations caused by external factors.
@@
Shallow intelligent models, such as support vector machines (Wen and Cao,, 2020; Wang et al.,, 2020) and decision trees (Cui et al.,, 2021), improve predictive accuracy by accommodating complex data patterns (Ahmad et al.,, 2014). Meanwhile, both feed-forward and recurrent neural networks have gained attention due to their ability to model nonlinearity and time-dependent structures within emission data (Wen and Yuan,, 2020; Ghalandari et al.,, 2021; Daniyal et al.,, 2022).
@@
The most recent innovations involve combined prediction models, which strategically integrate multiple individual models into a more robust and unified prediction framework (Tascikaraoglu and Uzunoglu,, 2014; Wang and Ye,, 2017). These hybrid approaches can be further categorised into combinations of statistical-statistical (Cui et al.,, 2018; Meng et al.,, 2014), statistical-intelligent (Wang et al.,, 2018; Zhao et al.,, 2018) and intelligent-intelligent (Acheampong and Boateng,, 2019; Li,, 2020). Each configuration offers distinct advantages in terms of accuracy, interpretability, and computational efficiency, making combined models a promising direction in emission forecasting.
@@ 1.2 Study contribution
Several model classes are explored. First, classical linear parametric models, including SARIMA and its extension with exogenous variables SARIMAX. Second, functional parametric models, such as the seasonal functional autoregressive (S)FAR and the functional autoregressive model with exogenous inputs (SFARX). Third, (semi) non-parametric, possibly non-linear, models, including generalized additive models (GAM) and the trigonometric seasonality, Box-Cox transformation, ARMA errors, trend, and seasonal (TBATS) modelling approach. Fourth, a "semi-functional" model is also considered as the K-nearest neighbours (KNN). All models are compared with a Naïve benchmark, in which the best prediction for a given hour is simply the value observed at the same hour the day before.
The comparison is further extended by evaluating forecast combination techniques that aggregate the predictions from the individual models described above. Specifically, combination methods include the simple average Bates and Granger, (1969), and a selection-based combination strategy.
Forecast performance is evaluated in terms of the hourly average root mean square error (RMSE), as well as the out-of-sample error $R^{2}$ and according to the Diebold-Mariano (DM) test.
The rest of the paper is structured as follows: CO2 emissions from electricity generation are described in Section 2; Section 3 introduces the single models (3.1) and forecast combination methods (3.2); the forecast experiment and its results are presented in Section 4; finally, Section 5 provides concluding remarks.
@@ 2 Data
Hourly carbon emissions from electricity generation, denoted as $E_{t}$, are calculated using the following emission factor-based method (Bertolini et al.,, 2025)
@@
Electricity generation data are sourced from the European Network of Transmission System Operators for Electricity (ENTSO-E,, 2025), and cover the 3-year period 1 January 2020 - 31 December 2023. Data are collected at both the national aggregate level and disaggregated across the seven electricity market zones: North, Centre-North, Centre-South, South, Calabria, Sicily and Sardinia. The data have an hourly frequency over 24 hours. The types of combusted fuels that generate electricity are: fossil brown coal, lignite, derived gas, natural gas, fossil hard coal, and fossil oil. Their country-specific emission factors and oxidation rates are sourced from the Italian institute for environmental protection and research (ISPRA,, 2025).
Figure 1 displays the CO2 emissions time series (in tonnes of CO2) from electricity generation in Italy over the period 1 January 2020 - 31 December 2023.

### v2

@@ Abstract
This study investigates the short-term forecasting of carbon emissions from electricity generation in the Italian power market. Using hourly data from 2021 to 2023, several statistical models and forecast combination methods are evaluated and compared at the national and zonal levels. Four main model classes are considered: (i) linear parametric models, such as seasonal autoregressive integrated moving average and its exogenous variable extension; (ii) functional parametric models, including seasonal functional autoregressive models, with and without exogenous variables; (iii) (semi) non-parametric and possibly non-linear models, notably the generalised additive model (GAM) and TBATS (trigonometric seasonality, Box-Cox transformation, ARMA errors, trend, and seasonality); and (iv) a semi-functional approach based on the K-nearest neighbours. Forecast combinations include simple averaging, the optimal Bates and Granger weighting scheme, and a selection-based strategy that chooses the best model for each hour. The results show that the GAM produces the most accurate forecasts during the daytime hours, while the functional parametric models perform best in the early morning. Among the combination methods, the simple average and the selection-based approaches consistently outperform all individual models. The findings underscore the value of hybrid forecasting frameworks in improving the accuracy and reliability of short-term carbon emissions predictions in power systems. In addition, they highlight the importance of considering zonal specificities when implementing flexible energy demand strategies, as the timing of low-carbon emissions varies between market zones throughout the day.
@@ 1 Introduction
Carbon dioxide (CO2) emissions are one of the main drivers of global climate change. According to the emissions database for global atmospheric research (edgar2024), in 2023, China, the world’s largest emitter, accounted for 33.98% of global CO2 emissions, followed by the USA (12%), India (7.57%), Europe (6.44%) and Russia (5.30%). The power sector remains a key contributor, responsible for 38.24% of total CO2 emissions in the same year. In particular, the power sectors of China, the USA, India, Europe and Russia contributed 16.59%, 3.75%, 3.53%, 2.26%, and 1.64% of global CO2 emissions, respectively. As a result, these countries represented 72.61% of global CO2 emissions from the power sector.
@@
The amount of CO2 emissions from electricity generation depends on the share of renewable sources (e.g. hydro, solar, and wind) and nonrenewable sources (e.g. coal, oil, and natural gas). The mix of sources changes over time according to the mechanisms of the electricity market and weather conditions (KOCAK2023). Although renewable energy generation reached a global record of 927 TWh in 2024, CO2 emissions from the power sector increased by 1.6% compared to 2023. This increase was mainly driven by an increase in global electricity demand, particularly due to extreme heat in countries such as India (carbonbrief2025). The rise in fossil fuel use highlights the ongoing challenge of decarbonising under climate stress.
@@
Over time, most countries have made efforts to reduce CO2 emissions through international initiatives. The Paris Agreement, signed in 2015, aims to keep global warming well below $2^{\circ}$C above pre-industrial levels, with the aim of staying below $1.5^{\circ}$C (UNFCCC2015). It also sets specific targets for the power sector, including a 20% reduction in CO2 emissions, a 20% increase in the use of renewable energy, and a 20% improvement in energy efficiency worldwide. These actions also support Sustainable Development Goal 7 (SDG 7), introduced in the same year with the 2030 Agenda for Sustainable Development, which focusses on ensuring access to affordable, reliable, and clean energy for all (SDG2015). Making the power sector more sustainable is essential to reduce CO2 emissions and address climate change.
@@
As part of the global shift towards cleaner energy in line with the climate targets of the United Nations (TARGET2050), more attention is paid to strategies that reduce CO2 emissions from electricity generation.
@@
From the supply side, proposed solutions include the integration of renewable energy sources with energy storage systems to manage fluctuations in renewable electricity generation. Batteries, fuel cells, or hydro reservoirs can be programmed to charge during periods of high renewable generation, typically associated with low emissions, and discharge when renewable output is low and consequently emissions are higher (HADDADIAN2015; IEA2019; He2024).
@@
On the demand side, there is growing interest in strategies that shift flexible energy demand from periods of high emissions to periods of low emissions. Smart charging represents an opportunity to reduce CO2 emissions by scheduling charging sessions during periods of lower emissions in electricity generation (JOCHEM2015). For example, residential charging of electric vehicles can be strategically planned when CO2 emissions are low (huber2018; HUBER2021). Even if electric vehicle drivers are not experts in the electric system (BIRESSELIOGLU2018), their willingness to adopt flexible charging, such as accepting delays, increases when it contributes to greater integration of renewable energy and reduces charging costs and carbon emissions (WILL2016; jung2019).
@@
Consumers typically schedule electricity usage based on price forecasts to minimise costs (voronin2013). However, price-based scheduling can lead to lower emissions only when renewable energy production exceeds the required (LEERBECK2020). Otherwise, it may result in a phenomenon known in the literature as the “merit order emission dilemma”, where low-cost but high-emission sources, such as coal, are preferred (wagner2002merit; FLESCHUTZ2021). Recently, more attention has been paid to the environmental impact of electricity consumption, with the aim of reducing CO2 emissions by promoting a shift to lower-carbon energy sources. In this context, electricity usage could be scheduled not only based on predicted prices but also on predicted CO2 emissions, as a primary or secondary objective (BOKDE2021). Therefore, accurate CO2 forecasts are essential to support emission-based scheduling of flexible demand and storage, at least until electricity generation becomes fully renewable.
Figure 1 shows the temporal structure of one day-ahead forecasts of carbon emissions from electricity generation in the day-ahead market. In this market, producers submit bids and offers for each hourly interval of day $t+1$. These transactions typically occur during the morning of day $t$, at which point the electricity generation for day $t+1$, as well as its associated carbon emissions, are unavailable. The red curve shows the fitted (forecasted) carbon emission profile for day $t+1$, while the black curve represents the true realisation observed ex-post. Accurate forecasts are crucial to identify low-carbon hours, namely periods during which electricity generation is expected to produce lower emissions. The grey highlighted intervals (00:00–06:00 and 11:00–13:00) indicate potential windows of demand flexibility, during which consumers may strategically shift electricity usage to optimise both economic outcomes (e.g., lower prices) and decarbonisation targets (BOKDE2021).
Figure 1: Time framework to forecast carbon emissions from electricity generation in the day-ahead market.
@@ 1.1 Literature review
In the CO2 forecasting framework, a distinction can be made between short-term and long-term CO2 emissions forecasting. Long-term forecasting, e.g. annual projections, enables policymakers to plan and manage future carbon footprints. By contrast, short-term CO2 emissions forecasting is essential for scheduling electricity production and consumption in ways that minimise resulting emissions over a day-ahead horizon. Although long-term forecasting has received substantial attention in the literature, studies focused on short-term forecasting are limited (BOKDE2021; LEERBECK2020). FINENKO2016 analyse hourly CO2 emissions from electricity generation in Singapore, highlighting hourly patterns due to both supply-side factors (e.g., renewable integration) and demand-side dynamics (e.g., peak vs. off-peak loads). MASON2018 applied evolutionary neural networks to forecast short-term CO2 emissions in Ireland over a multi-hour horizon. Lowry2018 forecasted the CO2 intensity of the United Kingdom’s electricity grid on a day-ahead horizon using seasonal autoregressive integrated moving average (SARIMA) models and artificial neural networks. LEERBECK2020 propose a machine learning algorithm to forecast short-term average and marginal CO2 emissions intensities in the Danish bidding zone DK2. The proposed algorithm combines multiple linear regression models through Softmax-weighted averaging, with residuals corrected using a SARIMA model with exogenous variables (SARIMAX). BOKDE2021 introduced two time series decomposition methods to forecast short-term CO2 emissions from electricity generation over the 48-hour horizon. The proposed methods are also compared with SARIMA models, model-based on differenced pattern sequences, feed-forward neural networks, pattern sequence-based forecasting methods, and support vector machine. Short-term CO2 emissions forecasts are obtained for five European countries: France, Germany, Poland, Denmark, and Norway. Maji2022b introduced the day-ahead carbon forecasting system, which forecasts the electricity generation of each source through an artificial neural network, then integrates these forecasts with the emission rates associated with each source to calculate the overall carbon emissions and carbon intensity. Maji2022a introduced CarbonCast, a machine learning system to forecast multi-day carbon intensities. CarbonCast has a hierarchical design that combine neural networks, weather forecasts, and historical data. Many of these studies (Lowry2018; LEERBECK2020; Maji2022b; ElectricityMap2022) provide forecasts for the day ahead, while some others provide multi-day forecasts (MASON2018; BOKDE2021; Maji2022a).
@@
Jin2024 provided a comprehensive review of carbon emission prediction models, identifying four main categories [^1] This classification closely aligns with that identified by Damon2002, who grouped models into three main categories: (i) time series and regression models-including ARMA models, threshold autoregressive models, generalized autoregressive conditional heteroskedasticity (GARCH) models, non-parametric regression models, generalized additive models, and adaptive regression splines; (ii) neural networks; and (iii) classification and regression tree models., [^2] Another interesting literature review on carbon emission forecast and model classification is provided by su17041471.: statistical models, shallow intelligent models, neural networks, and combined models.
@@
Statistical models are the most used because of their interpretability and reliance on historical data. Among them, the grey model is particularly popular for scenarios with limited data (deng1982). Regression models, such as linear and panel regression, have been used to analyse trade-related emissions and regional trends (SINGH2015271). Furthermore, time series models such as SARIMA and autoregressive distributed lag models (lin2019assessing) have proven effective in capturing temporal dynamics, offering good forecasting performance in various applications (LI2018337; Malik2020; Kaur2023). Despite their strengths in modelling trends and seasonality, these models may struggle with abrupt fluctuations caused by external factors.
@@
Shallow intelligent models, such as support vector machines (wen2020; Wang2020) and decision trees (cui2021), improve predictive accuracy by accommodating complex data patterns (AHMAD2014102). Meanwhile, both feed-forward and recurrent neural networks have gained attention due to their ability to model nonlinearity and time-dependent structures within emission data (wen20202; Ghalandari2021; daniyal2022).
@@
The most recent innovations involve combined prediction models, which strategically integrate multiple individual models into a more robust and unified prediction framework (TASCIKARAOGLU2014; WANG2017600). These hybrid approaches can be further categorised into combinations of statistical-statistical (cui2018; MENG2014673), statistical-intelligent (wang2018; zhao2018) and intelligent-intelligent (ACHEAMPONG2019; LiYan2020). Each configuration offers distinct advantages in terms of accuracy, interpretability, and computational efficiency, making combined models a promising direction in emission forecasting.
@@ 1.2 Study contribution
Several model classes are explored. First, classical linear parametric models SARIMAX. Second, functional parametric models, such as the seasonal functional autoregressive (S)FAR and the functional autoregressive model with exogenous inputs (SFARX). Third, (semi) non-parametric, possibly non-linear, models, including generalised additive models (GAM) and the trigonometric seasonality, Box-Cox transformation, ARMA errors, trend, and seasonal (TBATS) modelling approach. Fourth, a "semi-functional" model is also considered as the K-nearest neighbours (KNN). All models are compared with a Naïve benchmark, in which the best prediction for a given hour is simply the value observed at the same hour the day before.
The comparison is further extended by evaluating forecast combination techniques that aggregate the predictions from the individual models described above. Specifically, combination methods include the simple average, BG1969, and a selection-based combination strategy.
Forecast performance is evaluated using the hourly average root mean square error (RMSE), the out-of-sample $R^{2}$, the Diebold-Mariano (DM) test and the model confidence set (MCS).
The rest of the paper is structured as follows: CO2 emissions from electricity generation are described in Section 2; Section 3 introduces the single models (3.1) and forecast combination methods (3.2); the forecast experiment and its results are presented in Section 4; some implications of the policy on the scheduling of the electricity market are provided in Section 5; finally, Section 6 provides concluding remarks.
@@ 2 Data
Hourly carbon emissions from electricity generation, denoted as $E_{t}$, are calculated using the following emission factor-based method (BERTOLINI2025)
@@
Electricity generation data are sourced from the European Network of Transmission System Operators for Electricity (entsoe), and cover the 3-year period 1 January 2020 - 31 December 2023. Data are collected at both the national aggregate level and disaggregated across the seven electricity market zones: North, Centre-North, Centre-South, South, Calabria, Sicily and Sardinia. The data have an hourly frequency over 24 hours. The types of combusted fuels that generate electricity are: fossil brown coal, lignite, derived gas, natural gas, fossil hard coal, and fossil oil. Their country-specific emission factors and oxidation rates are sourced from the Italian institute for environmental protection and research (ispra).
Figure 2 displays the time series of CO2 emissions (in tonnes of CO2) from electricity generation in Italy over the period 1 January 2020 - 31 December 2023.
To address potential issues regarding structural breaks and instability, Figure 3 reports the rolling mean and standard deviation. These statistics are computed using two-week windows with a one-week overlap for the series of Italian carbon emissions (left panel) and their first differences (right panel) over the period 2021-2023. The results provide no indication of structural breaks, nor of instability in the mean or variance of the differenced series. A similar pattern is also observed across the individual market zones.
Figure 3: Two-week rolling mean and standard deviation computed with overlapping one-week windows, for the series of Italian carbon emissions (left panel) and its first differences (right panel) over the period 2021-2023.